OpenStudy (anonymous):

find the derivative of: y= ln[(3x-1)^3 (4x+2)^5)

OpenStudy (amistre64):

y = ln(u^3 v^5) u = 3x-1 ; v=4x+2 or y = ln(uv) ; u = (3x-1)^3 v=(4x+2)^5

OpenStudy (amistre64):

eiher way uv is on the bottom ans the top works out to be u'v' right?

OpenStudy (anonymous):

so i have to use the power rule?

OpenStudy (amistre64):

for the top yes; Dx(uv)

OpenStudy (anonymous):

thank yos so much. i just get confused when choosing which rule to use. =D

OpenStudy (amistre64):

uv' + u'v the top is: [(3x-1)^3] [20(4x+2)^4] + [9(3x-1)^2][(4x+2)^5] might wanna clean that up some tho...

OpenStudy (anonymous):

im going to work it out using the formula n then im going to check it up with your answer. thank you so much ^.^